Solve the following system by eliminating the y variable. (If there is no solution, enter NO SOLUTION for each answer. If there are infinitely many solutions, enter INFINITELY MANY for each answer.)
5x-3y=-7
15x-y=3
I need help. im really busting my brain with this one.

Answer by richwmiller(17219)

one solution
x = 2/5, y = 3
5x-3y=-7
multiply by 3
subtract
15x-9y=-21
15x-y=3
-8y=-24
y=3
5x-3(3)=-7
5x-9=-7
5x=2
x=2/5